# Using machine learning to measure evidence of students’ sensemaking in physics courses

### v2

@@ IV.1 ML Model Deployability Analysis
The best model for BoW achieved an average AUROC score of across validation folds with the hyper-parameters: . Similarly, for BERTfrozen, the best model achieved an AUROC score of with the hyper-parameters: . For BERTfinetuned, we require a larger amount of computational resources per experiment, so we only used a single training and validation set to run the hyperparameter grid search. Note that we use an L2SP loss [57, 58] when training BERTfinetuned, consisting of two separate weight decay terms such that we can regularize the weights from the BERT encoder separately from the weights of our LR classifiers. This is done to best preserve the prior information learned in in BERT’s original training. We set the batch-size to , which has been shown empirically to work well with deep neural networks [59]. Overall, we run 324 fine-tuning experiments for BERTfinetuned on Tufts’ HPC services. We picked the model with the optimal final AUROC of on the validation set, which had the following hyperparameters: , , and .
Model evaluations. We retrain our selected models on all training and validation data, and evaluate on the held-out test set of randomly selected students (60 student explanations). To approximate uncertainty on unseen student explanations, we bootstrap our test set with replacement to generate individual test sets for evaluating performance [60, 61]. In Figure 3, we show the average AUROC across independent test predictions for each of the 8 domain classifiers and for each model type. We are not able to report the AUROC for the properties domain as the class imbalance is extremely large and no positive examples were included in the test set. As such, we also provide the Area Under the Precision-Recall curve (AUPRC) results in Appendix E for all domain classifiers. This metric emphasizes the precision of the classifier, which informs us as to how well the model predicts the true positive labels. We note that even with AUPRC, there is too much uncertainty to make concrete claims about the deployability of the properties domain classifier. More positive data examples are necessary for improvement. We show the average AUROC across independent test set predictions for each model in Table 1. These values represent the annotation quality of our entire model. In Table 1, we also report the time taken to train the final selected model. To show the significance in results between models, we provide the confidence interval in Table 1 that results from our bootstrapping procedure and taking the AUROC differences between models. These values can be interpreted as the uncertainty intervals in AUROC difference across models, such that as long as zero is outside of the interval, we can be confident that there exists some non-zero difference between our models indicating that one is indeed better than the other. Additionally, we report the confusion matrices in Appendix F. BoW performs surprisingly well with an AUROC score of and trains efficiently in 1 minute. It performs nearly as well for objects as the other two models. This behavior is unsurprising as BoW reduces text to word frequency counts. BERTfrozen performs slightly better with an AUROC score of and takes 7 minutes to train. It performs the best for two of the domains: objects and mechanistic relationships. It is able to reliably outperform BoW on average, as 0 in not inside of the confidence interval between the AUROC classifier difference reported in Table 1. BERTfinetuned achieves an AUROC score of . It takes 160 minutes to train, but we emphasize that this is not a prohibitive cost for a model which only needs to be trained when new student data is available. It performs the best for influences, positioning, movements, interactions, and descriptive relationships. It is able to perform the best overall despite only leading in 5 domains due to the imbalance in the occurrence of each domain. We recommend the BERTfinetuned model as we believe the average increase in performance across domains is worth the nearly 3 hour training time, and the confidence intervals in 1 indicate it is a significant improvement over BERTfrozen. However, if the annotator wants to optimize for a particular domain, they should refer to Figure 3 to decide the best model for their purpose.
| AUROC (95% CI) | Runtime | |
@@
| 0.026 (0.005, 0.045) | |
| 0.077 (0.051, 0.104) | |
| 0.051 (0.024, 0.079) |
IV.2 Student Performance Analysis
We provide a quantitative snapshot of student performance in the first few weeks of the 2023 Tufts Intro. to Physics course. More specifically, we compute the sensemaking scores of students from their explanations to four problems (used in Sec. IV.1 and displayed in Appendix Sec. A), and compare these scores alongside problem-solving accuracies. Students reported their multiple-choice answer, such that accuracy here is a binary variable . Our results in Fig 4 show that assessing sensemaking scores provides a more nuanced view of student performance than correctness alone. In problem 1, we observe a high proportion of students solving the problem correctly () with very low sensemaking scores among correct answers (). In problem 2, we also observe high correctness (), but with higher sensemaking scores (an average of ) among correct answers. In these problems we observe students who obtain a correct answer with both high and low evidence of sensemaking. In fact, a closer look at the data reveals that many of these students who arrived at the correct answer without showing evidence of physical sensemaking deployed a learned mathematical rule (i.e. the derivative of velocity is acceleration).
In problems 3 and 4, we observe less correct answers across students ( and respectively); however sensemaking scores remain high among both correct (average sensemaking scores of and respectively) and incorrect ( and respectively) responses. Specifically in problem 4, the mean sensemaking scores for the incorrect group are higher than for the correct group. Recall that our measurement scheme determines evidence of the sensemaking process, not whether students have accurate information in the sensemaking process. Also, sensemaking has been characterized as a non-linear process to organize information that includes integrating ones phenomenological experiences, which often leads to constructive confusion [13, 14, 15, 16, 17]. Investigating the student responses reveals that students with higher sensemaking scores who arrived at the incorrect answer either used incorrect information or were unable to successfully make connections to their personal experiences with everyday physical objects.
| Problem | CI | |
| 1 | -0.166 | (-0.461, 0.145) |
| 2 | 0.053 | (-0.055, 0.176) |
| 3 | 0.404 | (0.227, 0.546) |
| 4 | -0.198 | (-0.391, 0.009) |
We show the Point-Biserial correlation coefficient , known as the binary-continuous form of the Pearson correlation coefficient ), for the student data in Table 2. The coefficient , where values farther from zero signals a stronger linear relationship and the sign indicates the direction. Only results in problem 3 suggest that there exists a positive linear correlation between problem-solving accuracy and sensemaking score; however, the coefficient is closer to zero than one indicating that the relationship strength is weak (). As we cannot assume normality in the underlying distributions over sensemaking scores for each binary group, we circumvent a hypothesis test with a bootstrapping procedure. We sample our data with replacement times and generate an empirical distribution over scores for each sample. We report the confidence interval in Table 2. All values within each interval remain closer to zero than , except for problem 3 for the tail end of the confidence interval. The weak correlation present makes a clear point: problem-solving accuracy is not a useful proxy for measuring evidence of physical sensemaking, nor vice versa. If one were to only look at the accuracy scores, one might believe that students are learning the core material. Sensemaking scores add more nuance to this belief, suggesting that students are engaging in alternative methods to obtaining a problem solution.
There are multiple actions one could take to further analyze these measurements and use them in a practical context, especially at the individual student level. For example, measurements could be used to observe whether evidence of sensemaking for a particular student is declining, stable, or growing across problems in time. To show an example of this, we show the trend for student in Fig. 4. This student shows evidence of stable sensemaking growth, which does not always coincide with correctness. In fact, if one only observed the accuracy score, the student’s performance would appear to be declining. When both of these measures are combined, one observes a more nuanced picture of the student’s progress, which could lead to more informed interventions.
V Conclusion & Future Work
In this work, we contribute a post-hoc measurement scheme to quantify evidence of physical sensemaking via human annotation of student explanations to physics problems. To streamline the lengthy process of annotating each student explanation, we automate the measurement scheme with a trained ML model using language encoders and shared probabilistic classifiers. We implement three model variants with different language encoders and conduct a deployability analysis on real student explanations from Tufts’ 2023 Intro. to Physics course. We show that the ML models have good average performance on data from new students (0.916 AUROC for the best model), but performance is dramatically different across domains. Moreover, we show that across problems, we can now systematically measure evidence of student learning separately from problem-solving accuracy. In fact, we observe no linear correlation between these two variables. Such a result has important implications for both practical instruction and STEM Ed. research. First and foremost it suggests that we need to be more cautious with how we measure learning. It’s not always useful to make assumptions about the process from the outcome. Students obtaining correct answers does not necessarily indicate a student has engaged in physical sensemaking, as there are alternative strategies, oftentimes requiring less time in the short-term, to achieve the same goal. As previous works support, sensemaking can more directly be measured by paying attention to how students solve problems - whether that’s through a post-hoc written explanation or an in-moment oral discussion. The former is where we’ve focused our efforts, but we believe the latter can also be more efficiently executed with ML tools. Beyond this, we perceive two next-stage research questions that lie at the intersection of ML and STEM Ed. fields:
(1) How can this post-hoc measurement scheme be improved? On the ML side, a natural first reaction is more data. We hypothesize that with more labeled training data across new problems, the models proposed here will be able to achieve better performance on the current task and generalize to unseen problems. Given collaborative annotation efforts, such a model could be realized more efficiently and would save all users time in the long run. We invite interested collaborators to get in touch and encourage open sharing of sensemaking data when possible and appropriate. One could also experiment with new state-of-the-art encoders to enhance overall performance [62]. For example, using the Mistral model (open access, open source transformer model in Ref. [63]) as the LLM text encoder, pre-trained with 7 billion parameters and 2 million language tokens. As off-the-shelf language models improve, one might not need as much training data to obtain quality performance. Lastly, our ML model assumes domain independence in the classification task. Perhaps there is a shared structure across domains that could be exploited. As a brief intuitive example, consider the perspective of an annotator. Given that the annotator identifies criteria in the first two domains (objects and influences), it is more probable that the annotator will identify criteria in the other domains. On the STEM Ed. side, we need to better understand how we can design problems that are aligned with the measurement scheme without creating a new performance metric for students. We do not want the annotation criterion to be perceived by students as tasks to complete. This runs the risk of students inventing alternative strategies to effectively mimic sensemaking. An example of this would be pattern recognizing the types of explanations that lead to high sensemaking scores. As shown by our ML model, this isn’t a very difficult task. Ideally, we prompt students to sensemake and our measurement scheme does not disturb their epistemic framing. There exists an open research pursuit in STEM Ed. to understand how problems can be effectively designed and communicated to students such that genuine sensemaking is encouraged [40].
(2) How can we model the dynamics that lead to low vs. high evidence of physical sensemaking in the classroom? Our proposed measurement scheme only provides evidence of a student’s sensemaking at a single point in time. We lack a model as to how students evolve such that specific sensemaking scores are realized. It is of high interest to construct both deterministic and probabilistic ML mechanistic models that capture short and long-term developmental dynamics of student sensemaking in physics classrooms[64, 65].
@@ Acknowledgements.
We are grateful for financial support from the U.S. National Science Foundation via awards GCR and CAREER . Furthermore, we wish to acknowledge Prof. David Hammer for providing us with the dataset necessary to carry out this work, high-level feedback on the draft, and thoughtful discussions regarding STEM Ed. literature. Without him, this work would not have been possible. We wish to acknowledge other members of the Hughes lab research group, namely Ethan Harvey and Patrick Feeney, for thoughtful ML discussions. Additionally, we’d like to thank the baristas at Jaho (Gabriel) and Starbucks (Ken) for supporting us through the final stages of this project.References
- Adams and Wieman [2015] W. K. Adams and C. E. Wieman, Analyzing the many skills involved in solving complex physics problems, American Journal of Physics 83, 459 (2015).
- Lestari et al. [2021] Lestari, S. Syafril, S. Latifah, E. Engkizar, D. Damri, Z. Asril, and N. E. Yaumas, Hybrid learning on problem-solving abiities in physics learning: A literature review, Journal of Physics: Conference Series 1796, 012021 (2021).
- Docktor et al. [2016] J. L. Docktor, J. Dornfeld, E. Frodermann, K. Heller, L. Hsu, K. A. Jackson, A. Mason, Q. X. Ryan, and J. Yang, Assessing student written problem solutions: A problem-solving rubric with application to introductory physics, Phys. Rev. Phys. Educ. Res. 12, 010130 (2016).
- Burkholder et al. [2020] E. W. Burkholder, J. K. Miles, T. J. Layden, K. D. Wang, A. V. Fritz, and C. E. Wieman, Template for teaching and assessment of problem solving in introductory physics, Phys. Rev. Phys. Educ. Res. 16, 010123 (2020).
- Morphew et al. [2018] J. W. Morphew, J. P. Mestre, H.-A. Kang, H.-H. Chang, and G. Fabry, Using computer adaptive testing to assess physics proficiency and improve exam performance in an introductory physics course, Phys. Rev. Phys. Educ. Res. 14, 020110 (2018).
- Lipnevich et al. [2020] A. A. Lipnevich, T. R. Guskey, D. M. Murano, and J. K. S. and, What do grades mean? variation in grading criteria in american college and university courses, Assessment in Education: Principles, Policy & Practice 27, 480 (2020).
- Fakcharoenphol et al. [2011] W. Fakcharoenphol, E. Potter, and T. Stelzer, What students learn when studying physics practice exam problems, Phys. Rev. ST Phys. Educ. Res. 7, 010107 (2011).
- Ben Ouahi et al. [2021] M. Ben Ouahi, M. Ait Hou, A. Bliya, T. Hassouni, and E. M. Al Ibrahmi, The effect of using computer simulation on students’ performance in teaching and learning physics: Are there any gender and area gaps?, Education Research International 2021, 6646017 (2021).
- Didi ş Körhasan and Pospiech [2024] N. Didi ş Körhasan and G. Pospiech, Examination of students’ quantum physics knowledge construction with peers and metacognitive evaluations, Phys. Rev. Phys. Educ. Res. 20, 020153 (2024).
- Laukenmann et al. [2003] M. Laukenmann, M. Bleicher, S. Fuß, M. Gläser-Zikuda, P. Mayring, and C. von Rhöneck, An investigation of the influence of emotional factors on learning in physics instruction, International Journal of Science Education 25, 489 (2003).
- Faridi et al. [2021] H. Faridi, N. Tuli, A. Mantri, G. Singh, and S. Gargrish, A framework utilizing augmented reality to improve critical thinking ability and learning gain of the students in physics, Computer Applications in Engineering Education 29, 258 (2021).
- Deslauriers et al. [2019] L. Deslauriers, L. S. McCarty, K. Miller, K. Callaghan, and G. Kestin, Measuring actual learning versus feeling of learning in response to being actively engaged in the classroom, Proceedings of the National Academy of Sciences 116, 19251 (2019).
@@
- Kaldaras and Wieman [2023a] L. Kaldaras and C. Wieman, Instructional model for teaching blended math-science sensemaking in undergraduate science, technology, engineering, and math courses using computer simulations, Phys. Rev. Phys. Educ. Res. 19, 020136 (2023a).
- Kaldaras and Wieman [2023b] L. Kaldaras and C. Wieman, Cognitive framework for blended mathematical sensemaking in science, Int. Journal of STEM Education https://doi.org/10.1186/s40594-023-00409-8 (2023b).
- KUO et al. [2013] E. KUO, M. M. HULL, A. GUPTA, and A. ELBY, How students blend conceptual and formal mathematical reasoning in solving physics problems, Science Education 97, 32 (2013).
- Bain et al. [2018] K. Bain, J.-M. G. Rodriguez, A. Moon, and M. H. Towns, The characterization of cognitive processes involved in chemical kinetics using a blended processing framework, Chem. Educ. Res. Pract. 19, 617 (2018).
- Bing and Redish [2007] T. J. Bing and E. F. Redish, The cognitive blending of mathematics and physics knowledge, AIP Conference Proceedings 883, 26 (2007).
- Gifford and Finkelstein [2020] J. D. Gifford and N. D. Finkelstein, Categorical framework for mathematical sense making in physics, Phys. Rev. Phys. Educ. Res. 16, 020121 (2020).
- Dreyfus et al. [2017] B. W. Dreyfus, A. Elby, A. Gupta, and E. R. Sohr, Mathematical sensemaking in quantum mechanics: An initial peek, Phys. Rev. Phys. Educ. Res. 13, 020141 (2017).
- Sherin [2001] B. L. Sherin, How students understand physics equations, Cognition and Instruction 19, 479 (2001).
- Note [1] Github repo: https://github.com/tufts-ml/AuSeM.git.
- Thagard and Kroon [2008] P. Thagard and F. Kroon, Hot Thought: Mechanisms and Applications of Emotional Cognition, A Bradford book (MIT Press, 2008).
@@
- Cronbach [1951] L. J. Cronbach, Coefficient alpha and the internal structure of tests., Psychometrik 16, 297–334 (1951).
- Amanda Benedict-Chambers and Palincsar [2017] E. A. D. Amanda Benedict-Chambers, Sylvie M. Kademian and A. S. Palincsar, Guiding students towards sensemaking: teacher questions focused on integrating scientific practices with science content, International Journal of Science Education 39, 1977 (2017).
@@
- Long et al. [2024] Y. Long, H. Luo, and Y. Zhang, Evaluating large language models in analysing classroom dialogue., npj Sci. Learn. (2024).
@@
- Hall and Krist [2025] K. Hall and C. Krist, Unsupervised ml with text data, in Applying Machine Learning in Science Education Research: When, How, and Why?, edited by P. Wulff, M. Kubsch, and C. Krist (Springer Nature Switzerland, Cham, 2025) pp. 281–312.

### v3

@@ IV RQ2: How ML Design Choices Can Impact Performance
Fine-tuning details. For BERTfinetuned, we require a larger amount of computational resources per experiment. Our grid search over hyperparameters entails 324 separate efforts to train this encoder’s millions of weights. We thus used a single training and validation set to run the hyperparameter grid search, avoiding the extra costs of 5-fold cross-validation. To better preserve the prior information learned in BERT’s original training, we use L2SP regularization [48, 49] when training BERTfinetuned. This means we enforce two separate sum-of-squared-difference penalties in the loss, one for BERT encoder weights and another for the LR classifier head weights. Each has a different desired mean and a different strength hyperparameter that is tuned. We set the batch-size to , which has been shown empirically to work well with deep neural networks [50].
During finetuning, we originally incorporated an domain classifier that we now believe to be out of scope for this particular work. As the domain-specific classifiers are independent, this does not influence the results of BoW or BERTfrozen. However, we acknowledge this may impact our reported performance of BERTfinetuned as the embeddings across classifiers are shared. We chose not to retrain as the model resources required to do another hyperparameter optimization (another set of 324 HPC experiments) is high. Re-training without the domain would likely only increase the performance of BERTfinetuned on 7 key domains, and as we show below it is already performing better than other variants .
Results from initial validation phase. The best model for BoW achieved an average AUROC score of across validation folds with the hyper-parameters: . Similarly, for BERTfrozen, the best model achieved an AUROC score of with the hyper-parameters: . For BERTfinetuned, after running 324 fine-tuning experiments on Tufts’ HPC services, we picked a model with final AUROC of on the validation set, which had the following hyperparameters: , , and .
Results on heldout test data. We retrain our selected models (with model-specific optimal hyperparameter settings) on all training and validation data. We then evaluate on the held-out test set of 60 responses from randomly selected students. In Fig. 3, we show the average AUROC across independent test predictions for each of the 7 domain classifiers and for each model type. To create approximate uncertainty intervals, we use bootstrap resampling of our test set [51, 52] to generate same-sized test sets that represent plausible unseen student responses, and compute the AUROC of each one. Intervals in Fig. 3 depict the 2.5-97.5th percentile range of these bootstrapped AUROCs.
Beyond the per-domain results in Fig. 3, we show the overall average AUROC for each model in Table 1. This micro-average is computed by weighting each annotation label equally across all domains. As such, these values represent the annotation quality of the multi-task classification set-up as a whole for each model. In Table 1, we also report the time taken to train the final selected model. To assess the significance of differences between models, we provide the confidence interval in Table 1 in the difference in AUROC between models, computed by taking percentiles from our bootstrapping procedure. These intervals can be interpreted as indicating uncertainty in AUROC difference across models. When comparing model A and B, if the interval is all positive and an exact zero is outside of the interval, we can be confident that there exists an actual positive difference between models, and model A is indeed better than B.
BoW performs decently overall with an AUROC score of and trains efficiently in 1 minute. For the objects domain in particular, it performs nearly as well as the other two models. BERTfrozen performs slightly better with an AUROC score of and takes 7 minutes to train. It performs the best for two domains: objects and chaining. It is able to reliably outperform BoW on average, as zero in not within the confidence interval for AUROC difference reported in Table 1. (The interval containing zero would indicate that no significant difference between models is plausible.) BERTfinetuned achieves an AUROC score of . It performs the best for influences, spatial organization, movement activities, and interactions. It takes significant time to train, 160 minutes. That time might not be prohibitive, depending on the resources available; training is only needed when new student data is available.
One major limitation in the results is that we cannot report AUROC results for yes/no criteria labels in the properties domain in Fig. 3. In our dataset, class imbalance is extremely large for this domain, and no positive examples were included in the test set, making AUROC undefined. More data, especially more positive examples, are necessary for improvement.
| AUROC (95% CI) | Train time | |
@@
| 0.029 (0.009, 0.049) | |
| 0.077 (0.052, 0.104) | |
| 0.048 (0.021, 0.075) |
V Benefits and Limitations
V.1 Utility.
As we described in the introduction, we are working to automate the measurement of mechanistic sensemaking evidence for research purposes across a set of student responses to conceptual problems in introductory Newtonian mechanics. This will allow larger-scale research studies than is possible with human annotating alone. Currently, annotating problems by hand is burdensome. Our annotator required about 5 hours to apply our measurement scheme to student responses for one problem. In contrast, once trained the processing time for our models to annotate the same amount of data is less than 1 minute. Furthermore, training a model to perform this task is doable in less than 3 hours. We hypothesize that with much more labeled training data across many additional problems, the model proposed here could perform well on unseen problem instances in introductory mechanics. However, the amount of data needed to cross such a threshold is unclear. Regardless, initial efforts to improve generalization across problems while reducing required annotation time should be directed towards collaborative annotation as well as semi-supervised learning methods [53].
Training models to measure mechanistic sensemaking on more data from new problems opens the door for interesting studies that have been difficult for PER. Consider this example hypothesis: “Evidence of mechanistic sensemaking early in a course is better than early canonical correctness at predicting both sensemaking and correct understanding at the end of the course.” Testing that hypothesis would involve many hours of human analysis; the ML we have begun to develop would be able to carry out this investigation quickly and efficiently. ML-enabled measurement could be used to conduct a randomized controlled experiment in PER, such as the one in Kuo et al. [54], but specifically for understanding the variables that lead to more mechanistic sensemaking in students. It could also be used to evaluate various introductory physics problems as being of low or high quality depending on whether a large enough body of students displayed (or did not display) evidence of mechanistic sensemaking in their responses (e.g. problem 1 in our dataset).
V.2 Generalizability
Russ et al. [22] developed a framework for assessing mechanistic reasoning that has been in use in PER for almost 20 years. To develop ML, we adapt the categories based on a particular dataset to produce a scheme with 7 domains. Researchers with similar datasets within the same conceptual domain of classical mechanics may find the scheme and the ML it supports useful as is. New contexts, however, may require new projections of Russ’s et al’s framework. For example, evidence of mechanistic reasoning in student explanations of electricity and magnetism or quantum mechanics will look different. It is likely that analogies would be an essential domain for some subject matter.
V.3 Reliability
By designing and defining the 7 domains for our dataset, and providing annotated examples, we hoped to reduce ambiguity for human annotators - those who might look to reproduce results on our data or to apply the scheme to their own. To measure this, we conducted an inter-rater reliability (IRR) test. Two annotators who took smaller roles in development of the domains, one professor of computer science and one senior undergraduate student in computer science, used the scheme to measure the last 15 student responses across the four physics problems (60 student responses total). The first author provided them with a codebook containing: the domain definitions, the four problems, the criteria vectors for each problem, and 5 annotated student examples per problem with some explicit reasons as to why certain coding decisions were made. Annotators provided initial ratings, then engaged in a collaborative discussion to resolve ambiguity and clarify disagreement. After this discussion, new ratings were assessed.
Among the three annotators (2 new raters alongside the first author who conducted all manual annotation for this work), the average accuracy and Cohen’s Kappa values [55] achieved pre-discussion are displayed in Table 3. The values that emerged post a discussion are presented in Table 3.
| Ave. Accuracy | Ave. Cohen’s Kappa | |
| Problem 1 | 1.0 | 1.0 |
| Problem 2 | 0.86 | 0.66 |
| Problem 3 | 0.82 | 0.63 |
| Problem 4 | 0.83 | 0.48 |
| Ave. Accuracy | Ave. Cohen’s Kappa | |
| Problem 1 | 1.0 | 1.0 |
| Problem 2 | 0.90 | 0.79 |
| Problem 3 | 0.94 | 0.88 |
| Problem 4 | 0.88 | 0.69 |
Problem 1 showed perfect agreement, but very few of the student examples show evidence of mechanistic sensemaking. (This suggests a form of utility for instruction: perhaps this is not a useful problem to assign.) For the remaining three problems, IRR scores were reasonably high ( accuracy). Cohen’s Kappa values incorporate chance due to data imbalance and are lower across problems, showing moderate to substantial agreement according to the categories in [55].
Discussion improved scores to accuracy, and all Cohen’s Kappa values increased to indicate substantial and almost perfect agreement [55]. After discussion, our codebook language was updated to improve its clarity.
For all results in Sec. IV, we trained ML to reproduce the labels provided by our single primary annotator. The 60 annotations from 2 other raters studied here were not included in our ML model training, as the IRR procedure was conducted in response to helpful reviewer feedback on the draft post model training based on the first author’s coding alone. As such, we highlight that this is a limitation of the current work - we do not know how our ML model will perform on measurements from multiple annotators nor do we have a model that corrects for mis-labeling. Investigating these directions is important for future larger-scale studies. Once an ML model is trained, the same student response will always receive the same annotation. Prediction is fully deterministic.
V.4 Sustainability Privacy.
As ML research advances, new encoders may be able to improve the overall classification performance. However, many of these models require High Performance Computing (HPC) access and large amounts of energy resources for training [56]. As language modeling in ML improves, it is worth benchmarking new methods while also taking into account the amount of compute resources required to obtain only modestly better performance. We only consider models which are open-access, which allows us to train classifiers for our specific task and does not require us to share student data with a 3rd party.
VI Conclusion & Future Work
In this work, we contribute a PER-ML measurement scheme to quantify evidence of mechanistic sensemaking in student responses to introductory mechanics problems. By introducing an ML implementation, our scheme can be more efficiently applied to many student responses on many problems than manual coding. This can enable large-scale quantitative research studies in PER, though this technology has both promises and limitations (see Sec. V) and there is no “magic bullet.”
In our work to date, we studied three model variants with different language encoders that require different levels of computational resources. We conducted a comparative analysis on real student responses. Results show that the ML models have good average performance on data from new students (0.916 AUROC for the best model). Performance does differ across domains of the scheme, in part reflecting the variance in the amount of available data per domain.
Throughout the research process, we identified new open questions and challenges. We thus conclude with reflections that we hope will motivate further research.
Recall we describe two steps of refinement from the prior framework of mechanistic reasoning [22]. The first was to design a scheme for measuring evidence of mechanistic reasoning in the context of student explanations in introductory mechanics. The second step was to train ML based on human annotated measurements for 100 students across four mechanics problems. Reflections below consider these steps in reverse order.
(1) How can the ML implementation of this measurement scheme be improved? A natural first reaction to this question is more data. We hypothesize that with more labeled training data across new problems, the supervised learning models proposed here will be able to achieve better performance on the current task and generalize to unseen problems. Given collaborative annotation efforts, such a model could be realized more efficiently and would save all users time in the long run. This also means that our model performance will depend on strong inter-rater agreement and accounting for human error in labeling. Considering future methods to account for different types of label noise is an interesting direction for future work.
There are a few paths that we could take to improve performance. The first is to experiment with newer encoders from the rapidly-advancing state-of-the-art frontier [57]. We could use the LLaMa encoder [2] previously used by Fussell et al. [8]. For an even newer encoder, we could explore the open-access, open-source 7-billion parameter Mistral model [35]. As off-the-shelf language models improve, one might not need as much task-specific training data to obtain quality performance, though it is hard to imagine doing well with less than our 4-problem, 100 student dataset. Task-specific training or in-context learning [58] will always be required. Out-of-the-box language encoders or encoder-decoder models (e.g. ChatGPT) are trained on general corpora of language, whereas the mechanistic sense-making of interest in PER reflects more particular disciplinary practices.
Another approach to improve the ML would be to leverage the dependency structure across our domains. As a brief intuitive example, consider the perspective of an annotator. If the annotator identifies that the student response did not discuss even the central object in the objects domain, it is essentially almost certain that properties, spatial organizations, and movement activities are also not provided. This correlation between domains could be exploited for better modeling, and may provide valuable inductive bias for achieving quality performance with less labeled data.
Lastly, we can pursue semi-supervised learning methods [53], which optimize model outputs to match labels that we do have, while also learning from patterns in a large, easier-to-collect un-labeled dataset (in our case, student response text without criteria vector annotations). We believe each of these paths forward are potentially valuable directions for improvement.
(2) How can new useful measurement schemes be achieved? We set out, and we remain interested, to assess student reasoning for sensemaking, in particular sensemaking relevant to the disciplinary practices of physics. For the present effort, we limited our attention to mechanistic sensemaking, drawing on prior work in PER and working to connect it to progress in ML. We limited our attention further to introductory mechanics. Other areas such as electricity and magnetism or quantum mechanics, would need different measurement schemes for evidence of mechanistic sensemaking. We are not prepared to hypothesize that ML could create such schemes, but it would be worth exploration. There are also many other approaches to explore for incorporating insight from PER and ML research. There are other bases in the literature regarding evidence of sensemaking, broader than mechanistic. There are also other data modalities (e.g. hand-written solutions with diagrams, graphs, and sketches; audio and visual recordings). These choices will influence our ML decisions, and vice versa. Depending on the ML constraints (e.g. complexity and interpretability of the model, amount of data and compute resources needed), frameworks from the PER literature will need to be adapted. Thus, as shown in this work, we put forth that the research process that integrates PER and ML is interactive. Ideas from the PER and ML side will influence one another until convergence - which is why we want to encourage collaborations between longtime members of the PER community and ML researchers who are interested in bringing their domain expertise to advancing the mission of PER. It is really through these collaborations that we can develop meaningful schemes that take advantage of the efficiency promise of ML, and unlock new insights in understanding student thinking.
@@ Acknowledgements.
We are grateful for financial support from the U.S. National Science Foundation via awards GCR and CAREER . We wish to acknowledge other members of the Hughes lab research group, namely Ethan Harvey and Cynthia Feeney, for thoughtful ML discussions. Additionally, we’d like to thank the baristas at Jaho (Gabriel) and Starbucks (Ken) for supporting us through the final stages of this project.References
- Touvron et al. [2023] H. Touvron, T. Lavril, G. Izacard, X. Martinet, M.-A. Lachaux, T. Lacroix, B. Rozière, N. Goyal, E. Hambro, F. Azhar, A. Rodriguez, A. Joulin, E. Grave, and G. Lample, Llama: Open and efficient foundation language models, arXiv preprint arXiv:2302.13971 (2023).
- BehnamGhader et al. [2024] P. BehnamGhader, V. Adlakha, M. Mosbach, D. Bahdanau, N. Chapados, and S. Reddy, Llm2vec: Large language models are secretly powerful text encoders (2024).
- Haab et al. [2022] J. Haab, N. Deutschmann, and M. R. Martínez, Is attention interpretation? a quantitative assessment on sets (2022), arXiv:2207.13018 [cs.LG] .
- Mylonas et al. [2022] N. Mylonas, I. Mollas, and G. Tsoumakas, An attention matrix for every decision: Faithfulness-based arbitration among multiple attention-based interpretations of transformers in text classification (2022), arXiv:2209.10876 [cs.CL] .
- Wulff et al. [2025] P. Wulff, M. Kubsch, and C. Krist, eds., Applying Machine Learning in Science Education Research: When, How, and Why?, 1st ed., Springer Texts in Education, Vol. 1 (Springer Cham, Cham, 2025) pp. XIII + 369, 19 b/w illustrations, 38 colour illustrations.
- Krist et al. [2025] C. Krist, M. Kubsch, and P. Wulff, Introduction, in Applying Machine Learning in Science Education Research: When, How, and Why?, edited by P. Wulff, M. Kubsch, and C. Krist (Springer Nature Switzerland, Cham, 2025) pp. 1–12.
- Fussell et al. [2025] R. K. Fussell, M. Flynn, A. Damle, M. F. J. Fox, and N. G. Holmes, Comparing large language models for supervised analysis of students’ lab notes, Phys. Rev. Phys. Educ. Res. 21, 010128 (2025).
@@
- Ko and Luna [2024] M.-L. M. Ko and M. J. Luna, The glue that makes it “hang together”: A framework for identifying how metadiscourse facilitates uncertainty navigation during knowledge building discussions, Journal of Research in Science Teaching 61, 457 (2024), https://onlinelibrary.wiley.com/doi/pdf/10.1002/tea.21893 .
- Phillips et al. [2017] A. M. Phillips, J. Watkins, and D. Hammer, Problematizing as a scientific endeavor, Phys. Rev. Phys. Educ. Res. 13, 020107 (2017).
- Machamer et al. [2000] P. Machamer, L. Darden, and C. F. Craver, Thinking about mechanisms, Philosophy of Science 67, 1 (2000).
- Watts et al. [2020] F. Watts, J. Schmidt-McCormack, C. Wilhelm, A. Karlin, A. Sattar, B. Thompson, A. Gere, and G. Shultz, What students write about when students write about mechanisms: analysis of features present in students’ written descriptions of an organic reaction mechanism, Chemistry Education Research and Practice 21 (2020).
@@
- Odden et al. [2020] T. O. B. Odden, A. Marin, and M. D. Caballero, Thematic analysis of 18 years of physics education research conference proceedings using natural language processing, Phys. Rev. Phys. Educ. Res. 16, 010142 (2020).
- Long et al. [2024a] Y. Long, H. Luo, and Y. Zhang, Evaluating large language models in analysing classroom dialogue., npj Sci. Learn. (2024a).
- Long et al. [2024b] Q. Long, Y. Wu, W. Wang, and S. J. Pan, Does In-Context Learning Really Learn? Rethinking How Large Language Models Respond and Solve Tasks via In-Context Learning, in First Conference on Language Modeling (2024).
- Min et al. [2022] S. Min, X. Lyu, A. Holtzman, M. Artetxe, M. Lewis, H. Hajishirzi, and L. Zettlemoyer, Rethinking the Role of Demonstrations: What Makes In-Context Learning Work?, in Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, edited by Y. Goldberg, Z. Kozareva, and Y. Zhang (Association for Computational Linguistics, Abu Dhabi, United Arab Emirates, 2022) pp. 11048–11064.
- Jiang et al. [2020] Z. Jiang, F. F. Xu, J. Araki, and G. Neubig, How Can We Know What Language Models Know?, Transactions of the Association for Computational Linguistics 8, 423 (2020).
- Liu et al. [2022] J. Liu, D. Shen, Y. Zhang, B. Dolan, L. Carin, and W. Chen, What Makes Good In-Context Examples for GPT-3?, in Proceedings of Deep Learning Inside Out (DeeLIO 2022): The 3rd Workshop on Knowledge Extraction and Integration for Deep Learning Architectures, edited by E. Agirre, M. Apidianaki, and I. Vulić (Association for Computational Linguistics, Dublin, Ireland and Online, 2022) pp. 100–114.
- Liang et al. [2021] P. P. Liang, C. Wu, L.-P. Morency, and R. Salakhutdinov, Towards Understanding and Mitigating Social Biases in Language Models, in Proceedings of the 38th International Conference on Machine Learning (PMLR, 2021) pp. 6565–6576.
@@
- van Engelen and Hoos [2020] J. E. van Engelen and H. H. Hoos, A survey on semi-supervised learning, Machine Learning 109, 373 (2020).
- Cohen [1960] J. Cohen, A coefficient of agreement for nominal scales, Educational and psychological measurement 20, 37 (1960).